The length of a rectangular playing field is 5 ft less than twice its width.
If the perimeter of the playing field is 230 ft, find the length and width of the field.

Answer by checkley75(3666)

2(2X-5)+2X=230
4X-10+2X=230
6X=230+10
6X=240
X=240/6
X=40 FOR THEWIDTH
2*40-5=80-5=75 FOR THE LENGTH.
PROOF
2*40+2*75=230
80+150=230
230=230